OpenStudy (anonymous):

write the quadratic equation whose roots are 2/3 and -1.

OpenStudy (anonymous):

(x-2/3)(x+1)=0 The problem can be rewritten as (3x-2)(x+1)=0

OpenStudy (a_clan):

let A and B be two roots. Then, quadratic equation is: x² - (A+B)x + AB = 0